# Unravelling in Collaborative Learning

### v1

@@ 1 Introduction
    We provide a rigorous framework for analyzing collaborative learning with strategic agents having data distributions of varying quality. On the one hand, we leverage tools from domain adaptation to capture a notion of data quality formally. On the other hand, collaboration is modeled as a principal-agent problem, where the principal is an aggregator in charge of collecting samples so as to maximize social welfare. This setup allows us to derive the benchmark welfare-maximizing collaboration scheme when data quality is public information.
@@ 2. Option 2)
Note that for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$, we take $n_{j}\geqslant 0$ to be a real number for ease of presentation. In (5), $a/c>0$ captures the trade-off between sampling burden and model quality.
@@
subject $\text{subject to}\quad\min_{j\in\mathopen{}\mathclose{{}\left[J}\right]}u_{j}((1,\mathsf{B}_{-j}),\mathbf{n})-\underline{u}_{j}\geqslant 0\;.$
@@ 3 Full-Information Benchmark: First-Best Collaboration
In this section, we assume that the profile of types $(\theta_{1},\ldots,\theta_{J})\in\Theta^{J}$ is public and study how the aggregator can implement an optimal collaboration among agents. For any $\mathsf{B}\in\{0,1\}^{J}$ and $\mathbf{n}\in\mathbb{R}_{+}^{J}$, denote by
@@
Working with the optimal scheme $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ is difficult because $\overline{n}_{j}({\bf 1},\mathbf{n}^{\mathrm{opt}}(\bm{\theta}))$ has no explicit expression. To make the analysis tractable, we consider the simplified optimal contribution scheme $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ where $\mathsf{B}^{\star}=\mathsf{B}^{\mathrm{opt}}=(1,\ldots,1)$ and for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$,
@@
Note that $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ only differs from $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ in two ways. First, in $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ all contributors’ participation constraint bind, while in $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ the $L^{\mathrm{opt}}$-th one could be slack. Second, the total number of data points required from the coalition is fixed and equal to $\bar{N}=\Theta(J^{\frac{1}{1+\gamma}})$. The alternative scheme $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ that we consider come from a natural relaxation of the original problem Equation 10 which is described in Appendix A.
The following proposition shows that $\mathbf{n}^{\star}(\bm{\theta})$ admits a workable expression.
@@
Applying $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ instead of $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ may come at a negligible welfare cost, as illustrated by the following result when types are sufficiently evenly spaced.
Lemma 2.
@@ 4.1 Naive aggregation and unravelling
A naive solution to coping with the private nature of $\bm{\theta}\in\Theta^{J}$ is for the aggregator to ask agents to disclose their types, and apply the contribution scheme defined in (12). In this setting, however, agents may declare a type $\tilde{\theta}_{j}$ different from their true type $\theta_{j}$.
@@
Theorem 2 shows that under $\Gamma$, the coalition undergoes a full unravelling: agent strategies result in a pure Nash equilibrium where the coalition is either empty or comprised solely of the worst agent. Thus, collaborative learning is not immune to adverse selection, and may suffer from unravelling as any market characterized by information asymmetry.
@@ 4.2 Breaking unravelling
Lemma 3 shows that some agent would need to receive a strictly positive transfer. This is impossible without a monetary payment (utility can only be decreased by the aggregator, for instance through accuracy shaping (Karimireddy et al., 2022)), which we exclude here.
@@
There exists a decreasing function $\eta_{\delta}:\>\mathbb{R}_{+}^{\star}\rightarrow\mathbb{R}_{+}^{\star}$, with $\delta\in(0,1)$ defined in H 1, such that for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$ and i.i.d samples $(X^{j}_{1},Y^{j}_{1}),\ldots,(X^{j}_{q},Y^{j}_{q})$ of size $q>0$ from $P_{j}$, there exists a measurable estimator $\widehat{\theta}_{j}$ of $(X^{j}_{1},Y^{j}_{1}),\ldots,(X^{j}_{q},Y^{j}_{q})$ satisfying

### v2

@@ 1 Introduction
    We provide a rigorous framework for analyzing collaborative learning with strategic agents having data distributions of varying quality. On the one hand, we leverage tools from domain adaptation to capture a notion of data quality formally. On the other hand, we model collaboration as a principal-agent problem, where the principal is an aggregator in charge of collecting samples so as to maximize social welfare. This setup allows us to derive the benchmark welfare-maximizing collaboration scheme when data quality is public information.
@@ 2. Option 2)
Note that for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$, we take $n_{j}\geqslant 0$ to be a real number for ease of presentation. In (5), $a/c>0$ captures the extent to which individuals are willing to trade off model quality against sampling cost.
@@
subject $\text{subject to}\quad\min_{j\in\mathopen{}\mathclose{{}\left[J}\right]}u_{j}(\mathsf{B},\mathbf{n})-\underline{u}_{j}\geqslant 0\;.$
@@ 3 Full-Information Benchmark: First-Best Collaboration
In this section, we assume that the profile of types $(\theta_{1},\ldots,\theta_{J})\in\Theta^{J}$ is public, and study how the aggregator can implement an optimal collaboration among agents under this most-favorable scenario.
@@
We are looking for a solution to the aggregator’s problem (8). For any $\mathsf{B}\in\{0,1\}^{J}$ and $\mathbf{n}\in\mathbb{R}_{+}^{J}$, denote by
@@
Working with the optimal scheme $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ is difficult because $\overline{n}_{j}({\bf 1},\mathbf{n}^{\mathrm{opt}}(\bm{\theta}))$ has no explicit expression. To make the analysis tractable, we slightly simply the optimal contribution scheme in Theorem 1 and consider the simplified optimal contribution scheme $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ where $\mathsf{B}^{\star}=\mathsf{B}^{\mathrm{opt}}=(1,\ldots,1)$ and for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$,
@@
Note that $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ only differs from $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ in two ways. First, in $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ all contributors’ participation constraint bind, while in $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ the $L^{\mathrm{opt}}$-the one could be slack. Second, the total number of data points required from the coalition is fixed and equal to $\bar{N}=\Theta(J^{\frac{1}{1+\gamma}})$. The quantity $\bar{N}$ comes from a natural relaxation of the original problem Equation 10 where we leave aside an intricate term of the objective function. This relaxation, which is formally described in Appendix A, provides a good approximation of the exact solution in reasonable settings. Indeed, the following result establishes that applying $(\mathsf{B}^{\star},\mathbf{n}^{\star})$ instead of $(\mathsf{B}^{\mathrm{opt}},\mathbf{n}^{\mathrm{opt}})$ comes at a negligible welfare cost when types are sufficiently evenly spaced.
Lemma 2.
Moreover, the following proposition shows that $\mathbf{n}^{\star}(\bm{\theta})$ admits a workable expression.
@@ 4.1 Naive aggregation and unravelling
A naive solution to coping with the private nature of $\bm{\theta}\in\Theta^{J}$ is for the aggregator to ask agents to disclose their types, and apply the simplified optimal contribution scheme defined in (12). In this setting, however, agents may declare a type $\tilde{\theta}_{j}$ different from their true type $\theta_{j}$.
@@
Theorem 2 shows that under $\Gamma$, the coalition undergoes a full unravelling: it is either empty or comprised solely of the worst agent in any Nash equilibrium. Thus, collaborative learning is not immune to adverse selection, and may suffer from unravelling as any market characterized by information asymmetry.
@@ 4.2 Breaking unravelling
Lemma 3 shows that some agent would need to receive a strictly positive transfer. This is impossible without a monetary payment–utility can only be decreased by the aggregator, for instance through accuracy shaping (Karimireddy et al., 2022)–, which we exclude here.
@@
There exists a decreasing function $\eta_{\delta}:\>\mathbb{R}_{+}^{\star}\rightarrow\mathbb{R}_{+}^{\star}$, with $\delta\in(0,1)$ defined in H 1, such that for any $j\in\mathopen{}\mathclose{{}\left[J}\right]$ and i.i.d samples $(X^{j}_{1},Y^{j}_{1}),\ldots,(X^{j}_{q},Y^{j}_{q})$ of size $q>0$ from $P_{j}$, there exists a $(X^{j}_{1},Y^{j}_{1}),\ldots,(X^{j}_{q},Y^{j}_{q})$-measurable estimator $\widehat{\theta}_{j}$ satisfying